# Fluctuations of Young diagrams for symplectic groups and semiclassical orthogonal polynomials

### v2

@@
Abstract.
Consider an $n\times k$ matrix of i.i.d. Bernoulli random numbers with $p=1/2$. Dual RSK algorithm gives a bijection of this matrix to a pair of Young tableaux of conjugate shape, which is manifestation of skew Howe $GL_{n}\times GL_{k}$-duality. Thus the probability measure on zero-ones matrix leads to the probability measure on Young diagrams proportional to the ratio of the dimension of $GL_{n}\times GL_{k}$-representation and the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{n}\otimes\mathbb{C}^{k}\right)$.
Similarly, by applying Proctor’s algorithm based on Berele’s modification of the Schensted insertion, we get skew Howe duality for the pairs of groups $Sp_{2n}\times Sp_{2k}$. In the limit when $n,k\to\infty$ $GL$-case is relatively easily studied by use of free-fermionic representation for the correlation kernel. But for the symplectic groups there is no convenient free-fermionic representation. We use Christoffel transformation to obtain the semiclassical orthogonal polynomials for $Sp_{2n}\times Sp_{2k}$ from Krawtchouk polynomials that describe $GL_{2n}\times GL_{2k}$ case. We derive an integral representation for semiclassical polynomials. The study of the asymptotic of this integral representation gives us the description of the limit shapes and fluctuations of the random Young diagrams for symplectic groups.
@@ Introduction
Consider an $n\times k$ matrix of i.i.d. Bernoulli random numbers with $p=1/2$. Dual Robinson-Schensted-Knuth algorithm gives a bijection of this matrix to a pair of semistandard Young tableaux of conjugate shape [1], which is manifestation of skew Howe $GL_{n}\times GL_{k}$-duality [2]. Thus the probability measure on zero-ones matrices leads to the probability measure on Young diagrams proportional to the ratio of the dimension of $GL_{n}\times GL_{k}$-representation and the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{n}\otimes\mathbb{C}^{k}\right)$. If $n$ and $k$ go to infinity with the same rate these random diagrams in French notation, rotated by $\frac{\pi}{4}$ and scaled by a factor $\frac{1}{n}$, converge uniformly in probability to a smooth limit shape supported on an interval. This limit shape and fluctuations around it were considered in [3, 4, 5].
@@
Random diagram can be seen as a random point configuration on a lattice, where a point is put in the left boundary of any unit interval, where the upper boundary of diagram decays. That is if diagram $\lambda$ has row lengths $(\lambda_{1},\dots,\lambda_{n})$ the coordinates of the points would be $\{\lambda_{i}+n-i\}_{i=1}^{n}$. Using Weyl dimension formula for irreducible representations of general linear groups it is possible to present this random point ensemble as orthogonal polynomial ensemble with Krawtchouk polynomials. The asymptotic behavior of Krawtchouk polynomials was studied extensively through the years, see [6, 7, 8, 9] and references therein. The corresponding determinatal ensemble was considered in [10, 11, 12, 13].
@@
Skew Howe duality for other pairs of classical Lie groups can be obtained by certain modifications of the insertion algorithm [14]. For example, by applying Proctor’s algorithm [14] based on Berele’s modification [15] of the Schensted insertion [16], we get skew Howe duality for the pairs of groups $Sp_{2n}\times Sp_{2k}$. The algorithm that proves $SO_{2n+1}\times\operatorname{Pin}_{2k}$ duality was proposed by Benkart and Stroomer [17], and for $O_{2n}\times SO_{2k}$ by Okada [18] and Terada [19]. All these algorithms give the bijections between zero-one matrices enumerating basis elements in an exterior algebra and pairs of corresponding modifications of semistandard Young tableaux for classical Lie groups, such as King tableaux for symplectic groups [20] and or Sundaram tableaux for orthogonal groups [21]. The duality leads to the multiplicity-free decomposition of the exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ into the direct sum of irreducible representations of dual groups pair. The decomposition gives rise to the probability measure on (symplectic or orthogonal) Young diagrams given by the ratio of dimension of the irreducible component to the dimension of the whole exterior algebra.
@@
Similarly to $GL_{n}\times GL_{k}$-case if we rotate and scale the diagram and take the limit $n,k\to\infty$ with the same rate, we get the uniform convergence to the limit shape. Moreover, the limiting curve is “half” of $GL_{2n}\times GL_{2k}$ curve and is described by the same explicit formula, as demonstrated in [22] by using the dimension formulas and solving the variational problems. The study of the fluctuations in these cases remained an open question.
@@
In the present paper we consider the orthogonal polynomial ensemble that corresponds to the $Sp_{2n}\times Sp_{2k}$-skew Howe duality. As demonstrated in [23], the polynomials are obtained from Krawtchouk polynomials by the “lifting” procedure from [24] which is a variant of the QR-algorithm for the Christoffel transformation of orthognal polynomials [25][26]. The polynomials are therefore discrete semiclassical orthogonal polynomials [27] and their asymptotic behavior was not considered previously. We study various asymptotic regimes for the orthogonal polynomials and obtain the description of local fluctuations for random symplectic Young diagrams.
For brevity we call these polynomials symplectic and they are new in this context. Our main results describe both the exact representation of the correlation kernel in terms of these polynomials and its asymptotic convergence in the bulk regime.
In what follows we introduce the notation:
  • Let $\{\widetilde{K}_{m}(x)\}_{m\geq 0}$ denote the system of monic Krawtchouk orthogonal polynomials on a suitable discrete set, orthogonal with respect to a weight function $W(x)$ (see (2.4)).
  • The modified weight is defined as
    $\widetilde{W}(x)=x^{2}\,W(x),$
    and the corresponding orthonormal system of symplectic polynomials (after applying the Christoffel transformation) is denoted by $\{g_{m}(x)\}_{m\geq 0}$.
  • The correlation kernel of the determinantal point process is given by the Christoffel–Darboux formula
    $\mathcal{K}(u,v)=2\frac{\varkappa_{2n-2}}{\varkappa_{2n}}\frac{g_{2n}(u)g_{2n-2}(v)-g_{2n-2}(u)g_{2n}(v)}{u^{2}-v^{2}}\sqrt{\widetilde{W}(u^{2})\widetilde{W}(v^{2})},$
    where $\varkappa_{m}$ is the leading coefficient of $g_{m}$, and the weight $\widetilde{W}(y)$ is expressed in terms of $y=x^{2}$.
We now state our main theorems.
Theorem 1 (Symplectic polynomials three term relation).
Define the lifted monic polynomials $\{G_{m}(x)\}$ by applying the Christoffel transformation (see (2.7)) corresponding to the change of weight $W(x)$ into $\widetilde{W}(x)=x^{2}\,W(x)$. The orthonormal symplectic polynomials are expressed as $g_{m}(x)=\frac{G_{m}(x)}{\sqrt{\Lambda_{m}}}$ with $\Lambda_{m}$ satisfying nonlinear recurrence relations (2.26),(2.27), (2.28). The coefficients in the three-term recurrence relation satisfied by the symplectic polynomials,
$x\,g_{k}(x)=\tilde{a}_{k}\,g_{k+1}(x)+\tilde{b}_{k}\,g_{k}(x)+\tilde{a}_{k-1}\,g_{k-1}(x),$
are obtained by applying the QR-algorithm to the recurrence coefficients of the original Krawtchouk polynomials (2.19). For $p=1/2$ the recurrence for $\Lambda_{m}$ simplifies to (2.46) and the recurrence for $\{\tilde{a}_{k},\tilde{b}_{k}\}$ is exactly solvable as (2.30) and (2.31). The kernel $\mathcal{K}(x,y)$ represents the orthogonal projection onto the span of $\{g_{0}(x),g_{2}(x),\dots,g_{2n-2}(x)\}$.
@@
Let $\mu_{n,k}$ be the measure corresponding to the $Sp_{2n}\times Sp_{2k}$ skew Howe duality, and let $\mathcal{K}(u,v)$ be the associated correlation kernel defined above. Define
@@
the kernel converges to the discrete sine kernel in the following sense:
@@
Theorem 1 is established by identifying the Christoffel transformation (2.7) as the mechanism to incorporate the weight modification (i.e., multiplication by $x^{2}$) and by analyzing the induced QR dynamics on the recurrence coefficients. Theorem 2 follows from a steepest descent analysis of the relevant integral representations for the polynomials and shows that the local fluctuations converge to universal sine kernel statistics; the density $\rho(x)$ is consistent with the limit shape analysis in [22].
@@
Acknowledgments
We thank Travis Scrimshaw for reading the first draft and pointing out the deficiencies in it. We thank Anton Dzhamay for useful discussions. The study was carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation in the framework of a scientific project under agreement No. 075-15-2025-013.
1. Random diagrams for symplectic groups
Consider the symplectic group $Sp_{2n}$ and its first fundamental representation on $\mathbb{C}^{2n}$. The exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\right)$ is a space of a reducible representation of $Sp_{2n}$. The exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ can be seen as $k$-th tensor power of this representation $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)=\left(\bigwedge\left(\mathbb{C}^{2n}\right)\right)^{\otimes k}$ and therefore has a natural action of $Sp_{2n}$. On the other hand the same can be done for $Sp_{2k}$, moreover, the actions of $Sp_{2n}$ and $Sp_{2k}$ commute on $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ and we have a strongly multiplicity-free decomposition into irreducible representations of $Sp_{2n}\times Sp_{2k}$ [28],[29]
@@ 1. Random diagrams for symplectic groups
Here $\lambda$ is a Young diagram that is confined inside of $n\times k$ rectangle, $\overline{\lambda}^{\prime}$ is its transposed complement in the rectangle and by $V_{Sp_{2n}}(\lambda)$ we denote the irreducible representation corresponding to the diagram $\lambda$. This is known as skew $(Sp_{2n},Sp_{2k})$ Howe duality [2], [30]. Combinatorially this duality can be understood in terms of King tableaux that generalize semistandard Young tableaux to the symplectic group case.
Let us briefly recall that King tableau [20] for the symplectic group $Sp_{2n}$ is Young diagram $\lambda$ filled in a semistandard way with signed numbers that have order $1<\bar{1}<2<\bar{2}<\dots<n<\bar{n}$. In English convention it means that numbers are strictly increasing along the columns and non-decreasing along the rows. An additional condition is imposed – all the numbers that are $\leq\bar{i}$ should not appear below row number $i$. King tableaux of shape $\lambda$ enumerate the basis vectors in the irreducible representation $V_{Sp_{2n}}(\lambda)$.
@@
The following algorithm based upon Berele insertion [15] was proposed by Proctor in [14] to establish the correspondence between zero-one matrices that encode the basis vectors in the left hand side of the decomposition (1.1) and pairs of King tableaux $(P,Q)$ of complementary shapes $(\lambda,\overline{\lambda}^{\prime})$ inside of $n\times k$ rectangle, that encode the basis vectors in the right hand side of the decomposition. We pair the columns of the zero-one matrix and on each step of the algorithm we consider the pair of columns. On step $r$ the shapes of $P,Q$ form the $n\times r$ rectangle. From two consecutive columns number $12r,2r+$ 1 we produce a sequence of numbers to insert by reading the columns bottom-up. If in row number $i$ we have $(0,0)$ we produce the number $i$, for $(0,1)$ we insert nothing, for $(1,0)$ we insert $\bar{i},i$ and for $(1,1)$ we insert $\bar{i}$. Then we insert numbers one-by-one by using Schensted insertion [16]. If the additional condition is broken by an insertion of number $i$ that is bumping number $\bar{i}$ below the row $i$, we need to erase the box where $\bar{i}$ was. Then we shift this empty box to the boundary as in jeu de taquin (to preserve the semi-standard condition). After the insertion of all the numbers at step $r$, the erased boxes are filled with numbers $\bar{r}$ and adjoined to $Q$. All the empty boxes between the extended tableaux $P$ and $Q$ in $n\times(r+1)$-rectangle are filled with $r$ and adjoined to $Q$ as well. Then we proceed to the step number $(r+1)$.
@@
It is possible to write this measure explicitly, as was demonstrated in [22]. To do so introduce the coordinates $a_{i}=\lambda_{i}+n-i+1$. These coordinates can be interpreted as the coordinates of unit intervals where the upper boundary of Young diagram $\lambda$, rotated by $\frac{\pi}{4}$ (“Russian convention”), is decreasing. The diagram then can be identified with a particle configuration $\left\{a_{i}\right\}$ on the integer lattice. For $\dim V_{Sp_{2n}}(\lambda)$ we can use Weyl dimension formula. The problem is to write the dimension of $V_{Sp_{2k}}(\overline{\lambda}^{\prime})$ in terms of the coordinates $\{a_{i}\}$, which are introduced for the diagram $\lambda$. It can be done by representing King tableau as a set of non-intersecting paths, then switching to the dual paths and using Lindström–Gessel–Viennot lemma to express the number of non-intersecting paths configurations as a determinant [31]. Then Dodgson condensation leads to the explicit formula
$\mu_{n,k}(\left\{a_{i}\right\})=\frac{2^{2n(1-k)}}{\prod_{i<j}(j-i)(2n+2-i-j)}\prod_{i<j}(a_{i}^{2}-a_{j}^{2})^{2}\prod_{\ell=1}^{n}a_{\ell}^{2}\frac{(2k-1+2\ell)!}{(k+n+a_{\ell})!(k+n-a_{\ell})!}.$ (1.3)
By this derivation the measure is determinantal, therefore the measure and all of the correlation functions can be expressed as the determinants of the correlation kernel
$\mu_{n,k}(\left\{a_{i}\right\})=\det\left[\mathcal{K}(a_{i},a_{j})\right]_{i,j=1}^{n}.$ (1.4)
@@
In [22] we have demonstrated that random Young diagrams, rotated by $\frac{\pi}{4}$ and considered as 1-Lipschitz functions, converge uniformly in probability to the limit curve $\Omega(u)$ that has an explicit formula for the limit density $\mathbb{P}(\sup_{u}|\lambda(u)-\Omega(u)|>\varepsilon)\to 0$ as $n\to\infty$. In [23] the same was demonstrated for a very similar problem in $(SO_{2n+1},\operatorname{Pin}_{2k})$ skew Howe duality. To do so we have written the measure in the exponential form, used Stirling approximation for the factorials and solved the variational problem. But we have not considered the local fluctuations around the limit shape. Nevertheless in [23] in a similar problem we have demonstrated that the probability measure can be written in terms of a semiclassical family of orthogonal polynomials. This representation was then used to prove Central Limit Theorem for the global fluctuations. Below we use the polynomials to write the correlation kernel in the Christoffel-Darboux form and demonstrate its convergence to the discrete sine kernel in the bulk by analyzing the asymptotic behavior of the orthogonal polynomials.
@@
$Z_{n,k}^{-1}=\frac{2^{2n(1-k)}n^{2n^{2}}}{\prod\limits_{i<j}^{n}(j-i)(2n+2-i-j)}\prod\limits_{l=1}^{n}\frac{(2k-1+2l)!}{(2k+2n)!}$
@@
According to (1.7), the weight $\widetilde{W}(u^{2})$ is the binomial coefficient multiplied by $u^{2}$. The polynomials, that are orthogonal with the binomial weight are the Krawtchouk polynomials. The polynomials $p_{m}$ are related to the Krawtchouk polynomials using the “lifting” procedure from [24] which is a variant of the QR-algorithm for the Christoffel transformation of orthognal polynomials [25][26] as demonstrated below.
@@
$\mu_{n,k}(y_{1},\dots,y_{n})=\frac{1}{Z_{n,k}}\det\left[\mathcal{K}(y_{i},y_{j})\right]_{i,j=1}^{n}=\frac{1}{Z_{n,k}}\det\left(\sqrt{\widetilde{W}(y_{i})\widetilde{W}(y_{j})}\sum_{l=0}^{n-1}p_{l}(y_{i})\;p_{l}(y_{j})\right)_{i,j=1}^{n}.$ (1.10)
@@
2. A family of semiclassical orthogonal polynomials
2.1. Symplectic polynomials

### v3

@@
Abstract
Consider an $n\times k$ matrix of i.i.d. Bernoulli random numbers with $p=1/2$. The dual RSK algorithm gives a bijection of this matrix to a pair of Young tableaux of conjugate shape, which is manifestation of skew Howe $GL_{n}\times GL_{k}$-duality. Thus the probability measure on zero-ones matrix leads to the probability measure on Young diagrams proportional to the ratio of the dimension of $GL_{n}\times GL_{k}$-representation and the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{n}\otimes\mathbb{C}^{k}\right)$.
Similarly, by applying Proctor’s algorithm based on Berele’s modification of the Schensted insertion, we get skew Howe duality for the pairs of groups $Sp_{2n}\times Sp_{2k}$. In the limit when $n,k\to\infty$ GLGL-case is relatively easily studied by use of free-fermionic representation for the correlation kernel. But for the symplectic groups there is no convenient free-fermionic representation. We use Christoffel transformation to obtain the semiclassical orthogonal polynomials for $Sp_{2n}\times Sp_{2k}$ from Krawtchouk polynomials that describe $GL_{2n}\times GL_{2k}$ case. We derive an integral representation for semiclassical polynomials. The study of the asymptotic of this integral representation gives us the description of the limit shapes and fluctuations of the random Young diagrams for symplectic groups.
keywords:
Young diagrams, Symplectic groups, Skew Howe duality, Semiclassical orthogonal polynomials, Christoffel transformation, Krawtchouk polynomials, Correlation kernel and sine kernel limit, Asymptotic analysis, limit shape and fluctuations
pacs:
MSC Classification]33C45, 60G55, 22E46, 60B10
@@ Introduction
Consider an $n\times k$ matrix of i.i.d. Bernoulli random numbers with $p=1/2$. Dual Robinson-Schensted-Knuth algorithm gives a bijection of this matrix to a pair of semistandard Young tableaux of conjugate shape [knuth1970permutations], which is manifestation of skew Howe $GL_{n}\times GL_{k}$-duality [R_Howe]. Thus the probability measure on zero-ones matrices leads to the probability measure on Young diagrams proportional to the ratio of the dimension of $GL_{n}\times GL_{k}$-representation and the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{n}\otimes\mathbb{C}^{k}\right)$. If $n$ and $k$ go to infinity with the same rate these random diagrams in French notation, rotated by $\frac{\pi}{4}$ and scaled by a factor $\frac{1}{n}$, converge uniformly in probability to a smooth limit shape supported on an interval. This limit shape and fluctuations around it were considered in [GTW01, GTW02, GTW02II].
@@
Random diagram can be seen as a random point configuration on a lattice, where a point is put in the left boundary of any unit interval, where the upper boundary of diagram decays. That is if diagram $\lambda$ has row lengths $(\lambda_{1},\dots,\lambda_{n})$ the coordinates of the points would be $\{\lambda_{i}+n-i\}_{i=1}^{n}$. Using Weyl dimension formula for irreducible representations of general linear groups it is possible to present this random point ensemble as orthogonal polynomial ensemble with Krawtchouk polynomials. The asymptotic behavior of Krawtchouk polynomials was studied extensively through the years, see [ISMAIL1998121, BKMM03, LI2000155, Wang_Qui_2015] and references therein. The corresponding determinatal ensemble was considered in [johansson2002non, Borodin_2003, borodin2007asymptotics, Borodin_2017].
@@
Skew Howe duality for other pairs of classical Lie groups can be obtained by certain modifications of the insertion algorithm [proctor1993reflection]. For example, by applying Proctor’s algorithm [proctor1993reflection] based on Berele’s modification [berele1986schensted] of the Schensted insertion [schensted1961longest], we get skew Howe duality for the pairs of groups $Sp_{2n}\times Sp_{2k}$. The algorithm that proves $SO_{2n+1}\times\operatorname{Pin}_{2k}$ duality was proposed by Benkart and Stroomer [benkart1991tableaux], and for $O_{2n}\times SO_{2k}$ by Okada [okada1993robinson] and Terada [Terada93]. All these algorithms give the bijections between zero-one matrices enumerating basis elements in an exterior algebra and pairs of corresponding modifications of semistandard Young tableaux for classical Lie groups, such as King tableaux for symplectic groups [King76] and or Sundaram tableaux for orthogonal groups [Sundaram90]. The dualities lead to the multiplicity-free decompositions of the exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ into the direct sums of irreducible representations of dual groups pairs. The decompositions give rise to the probability measures on (symplectic or orthogonal) Young diagrams $\lambda$ given by the ratio of dimension of the irreducible component to the dimension of the whole exterior algebra. It is possible to write these measures explicitly, as was demonstrated in [nazarov2024skew].
@@
Similarly to $GL_{n}\times GL_{k}$-case if we rotate and scale the diagram and take the limit $n,k\to\infty$ with the same rate, we get the uniform convergence to the limit shape. Moreover, the limiting curve is “half” of $GL_{2n}\times GL_{2k}$ curve and is described by the same explicit formula, as demonstrated in [nazarov2024skew] by using the dimension formulas and solving the variational problems. The study of the fluctuations in these cases remained an open question.
In the present paper we consider the $Sp_{2n}\times Sp_{2k}$-skew Howe duality. To write the measure explicitly introduce the coordinates $a_{i}=\lambda_{i}+n-i+1$, then as explained in the Section 1, we get
$\mu_{n,k}(\left\{a_{i}\right\})=\frac{1}{Z_{n,k}}\prod_{i<j}(a_{i}^{2}-a_{j}^{2})^{2}\prod_{\ell=1}^{n}a_{\ell}^{2}\binom{2n+2k}{n+k+a_{\ell}}=\frac{1}{Z_{n,k}}\det\left[\mathcal{K}\left(\frac{a_{i}}{n},\frac{a_{j}}{n}\right)\right]_{i,j=1}^{n},$ (0.1)
where $Z_{n,k}$ is the normalization constant, given in (1.7). The measure contains a Vandermonde determinant written in terms of squares of variables $a_{i}$. Therefore it can be expressed as the determinant of the Cristoffel-Darboux kernel $\mathcal{K}\left(u,v\right)$ for certain orthogonal polynomials. For brevity we call these polynomials symplectic and they are new in this context.
@@
As demonstrated in [NNP20], symplectic polynomials are obtained from Krawtchouk polynomials by the “lifting” procedure from [BUHMANN1992117] which is a variant of the QR-algorithm for the Christoffel transformation of orthognal polynomials [galant1971implementation][bueno2004darboux]. The polynomials are therefore discrete semiclassical orthogonal polynomials [van_assche_2017] and their asymptotic behavior was not considered previously. We study various asymptotic regimes for the orthogonal polynomials and obtain the description of local fluctuations for random symplectic Young diagrams in the bulk regime, given by the next theorem.
@@
Let $\mu_{n,k}$ be the measure corresponding to the $Sp_{2n}\times Sp_{2k}$ skew Howe duality (0.1), and let $\mathcal{K}(u,v)$ be the associated correlation kernel defined above. Define
@@
the kernel converges pointwisely to the discrete sine kernel:
@@
Theorem 1 follows from a steepest descent analysis of the relevant integral representations for the polynomials and shows that the local fluctuations converge to universal sine kernel statistics; the density $\rho(x)$ is consistent with the limit shape analysis in [nazarov2024skew].
@@
1 Random diagrams for symplectic groups
Consider the symplectic group $Sp_{2n}$ and its first fundamental representation on $\mathbb{C}^{2n}$. The exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\right)$ is a space of a reducible representation of $Sp_{2n}$. The exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ can be seen as $k$-th tensor power of this representation $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)=\left(\bigwedge\left(\mathbb{C}^{2n}\right)\right)^{\otimes k}$ and therefore has a natural action of $Sp_{2n}$. On the other hand the same can be done for $Sp_{2k}$, moreover, the actions of $Sp_{2n}$ and $Sp_{2k}$ commute on $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ and we have a strongly multiplicity-free decomposition into irreducible representations of $Sp_{2n}\times Sp_{2k}$ [cheng2012dualities],[goodman2009symmetry]
@@ 1 Random diagrams for symplectic groups
Here $\lambda$ is a Young diagram that is confined inside of $n\times k$ rectangle, $\overline{\lambda}^{\prime}$ is its transposed complement in the rectangle and by $V_{Sp_{2n}}(\lambda)$ we denote the irreducible representation corresponding to the diagram $\lambda$. This is known as skew $(Sp_{2n},Sp_{2k})$ Howe duality [R_Howe], [howe1995perspectives]. Combinatorially this duality can be understood in terms of King tableaux that generalize semistandard Young tableaux to the symplectic group case.
Let us briefly recall that King tableau [King76] for the symplectic group $Sp_{2n}$ is Young diagram $\lambda$ filled in a semistandard way with signed numbers that have order $1<\bar{1}<2<\bar{2}<\dots<n<\bar{n}$. In English convention it means that numbers are strictly increasing along the columns and non-decreasing along the rows. An additional condition is imposed – all the numbers that are $\leq\bar{i}$ should not appear below row number $i$. King tableaux of shape $\lambda$ enumerate the basis vectors in the irreducible representation $V_{Sp_{2n}}(\lambda)$.
@@
The following algorithm based upon Berele insertion [berele1986schensted] was proposed by Proctor in [proctor1993reflection]. It establishes the correspondence between zero-one matrices of size $n\times 2k$, that encode the basis vectors in the left hand side of the decomposition (1.1), and pairs of King tableaux $(P,Q)$ of complementary shapes $(\lambda,\overline{\lambda}^{\prime})$ inside of $n\times k$ rectangle, that encode the basis vectors in the right hand side of the decomposition.
  • First, we pair the columns of the zero-one matrix. On each step of the algorithm we consider the pair of columns, so total number of algorithm steps is $k$.
  • We start with the empty tableaux $P$ and $Q$. On the first step we fill the $n\times 1$ rectangle with numbers and adjoin some of the boxes to the tableau $P$ and remaining to the tableau $Q$ as explained below.
  • After the step r1r-1 the shapes of $P,Q$ form the $n\times(r-1)$ rectangle. We start the step $r$ by shifting the tableau $Q$ to the right by one box, so that there appears an empty strip between $P$ and $Q$ in $n\times r$ rectangle.
  • From two consecutive columns number $2r,2r+1$ we produce a sequence of numbers to insert on the step $r$ by reading the columns bottom-up as follows.
  • If in row number $i$ we have $(0,0)$ in the columns, we produce the number $i$, for $(0,1)$ we insert nothing, for $(1,0)$ we insert $\bar{i},i$ and for $(1,1)$ we insert $\bar{i}$.
  • Then we insert numbers one-by-one by using Schensted insertion [schensted1961longest] into tableau $P$. If the additional condition is broken by an insertion of number $i$ that is bumping number $\bar{i}$ below the row $i$, we need to erase the box where $\bar{i}$ was. Then we shift this erased box to the boundary as in jeu de taquin (to preserve the semi-standard condition).
  • After the insertion of all the numbers at step $r$, the erased boxes are filled with numbers $\bar{r}$ and adjoined to $Q$.
  • All the remaining empty boxes between the extended tableaux $P$ and $Q$ in $n\times r$-rectangle are filled with $r$ and adjoined to $Q$ as well. Then we proceed to the step number $(r+1)$.
@@
It is possible to write this measure explicitly, as was demonstrated in [nazarov2024skew]. To do so introduce the coordinates $a_{i}=\lambda_{i}+n-i+1$. These coordinates can be interpreted as the coordinates of unit intervals where the upper boundary of Young diagram $\lambda$, rotated by $\frac{\pi}{4}$ (“Russian convention”), is decreasing. The diagram then can be identified with a particle configuration $\left\{a_{i}\right\}$ on the integer lattice. For $\dim V_{Sp_{2n}}(\lambda)$ we can use Weyl dimension formula. The problem is to write the dimension of $V_{Sp_{2k}}(\overline{\lambda}^{\prime})$ in terms of the coordinates $\{a_{i}\}$, which are introduced for the diagram $\lambda$. It can be done by representing King tableau as a set of non-intersecting paths, then switching to the dual paths and using Lindström–Gessel–Viennot lemma to express the number of non-intersecting paths configurations as a determinant [oh2019identities]. Then Dodgson condensation leads to the explicit formula
$\mu_{n,k}(\left\{a_{i}\right\})=$
$=\frac{2^{2n(1-k)}}{\prod_{i<j}(j-i)(2n+2-i-j)}\prod_{i<j}(a_{i}^{2}-a_{j}^{2})^{2}\prod_{\ell=1}^{n}a_{\ell}^{2}\frac{(2k-1+2\ell)!}{(k+n+a_{\ell})!(k+n-a_{\ell})!}.$ (1.3)
From this derivation the measure is determinantal, and the associated point process is therefore a determinantal point process (see [borodin2009determinantalpointprocesses] and references therein). Consequently, the measure can be written as determinant of the correlation kernel,
$\mu_{n,k}(\{a_{i}\})=\det\big[\mathcal{K}(a_{i},a_{j})\big]_{i,j=1}^{n}.$ (1.4)
@@
In [nazarov2024skew] we have demonstrated that random Young diagrams, rotated by $\frac{\pi}{4}$ and considered as 1-Lipschitz functions, converge uniformly in probability to the limit curve $\Omega(u)$ that has an explicit formula for the limit density $\mathbb{P}(\sup_{u}|\lambda(u)-\Omega(u)|>\varepsilon)\to 0$ as $n\to\infty$. In [NNP20] the same was demonstrated for a very similar problem in $(SO_{2n+1},\operatorname{Pin}_{2k})$ skew Howe duality. To do so we have written the measure in the exponential form, used Stirling approximation for the factorials and solved the variational problem. But we have not considered the local fluctuations around the limit shape. Nevertheless in [NNP20] in a similar problem we have demonstrated that the probability measure can be written in terms of a semiclassical family of orthogonal polynomials. This representation was then used to prove Central Limit Theorem for the global fluctuations. Below we use the polynomials to write the correlation kernel in the Christoffel-Darboux form and demonstrate its convergence to the discrete sine kernel in the bulk by analyzing the asymptotic behavior of the orthogonal polynomials.
@@
$Z_{n,k}^{-1}=\frac{2^{2n(1-k)}n^{2n^{2}}}{\prod\limits_{i<j}^{n}(j-i)(2n+2-i-j)}\prod\limits_{l=1}^{n}\frac{(2k-1+2l)!}{(2k+2n)!}$ (1.7)
@@
According to (1.8), the weight $\widetilde{W}(u^{2})$ is the binomial coefficient multiplied by $u^{2}$. The polynomials, that are orthogonal with the binomial weight are the Krawtchouk polynomials. The polynomials $p_{m}$ are related to the Krawtchouk polynomials using the “lifting” procedure from [BUHMANN1992117] which is a variant of the QR-algorithm for the Christoffel transformation of orthognal polynomials [galant1971implementation][bueno2004darboux] as demonstrated below.
@@
$\mu_{n,k}(y_{1},\dots,y_{n})=$ $\frac{1}{Z_{n,k}}\det\left[\mathcal{K}(y_{i},y_{j})\right]_{i,j=1}^{n}=$
$=$ $\frac{1}{Z_{n,k}}\det\left(\sqrt{\widetilde{W}(y_{i})\widetilde{W}(y_{j})}\sum_{l=0}^{n-1}p_{l}(y_{i})\;p_{l}(y_{j})\right)_{i,j=1}^{n}.$ (1.11)
@@
2 A family of semiclassical orthogonal polynomials
2.1 Symplectic polynomials